study the converges unformly ?
* au f (x) x W xE68] study the roaware unformlyz Pux). aut tx Głos) * fn U l = X en Neolus)
please provide detailed step by step solution
Real analysis
Question in Real analysis
Answer #1

n-> #n>NEN Dhene eso řs arbitary tue fon). 3 ffx) = x21 n6 [,87 . How lim fn(x) = lim n = 0 = f(x) The point wise limit of fa© , x) = xe-n* 4- x = lẽ, 8 Point wise limit of frin) is given by lingo hen) = nuimme on = 0 = fen). a fon Exo, MEN sit [fn (